# Full Parameter Fine-tuning for Large Language Models with Limited Resources

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) but demand massive GPU resources for training. Lowering the threshold for LLMs training would encourage greater participation from researchers, benefiting both academia and society. While existing approaches have focused on parameter-efficient fine-tuning, which tunes or adds a small number of parameters, few have addressed the challenge of tuning the full parameters of LLMs with limited resources. In this work, we propose a new optimizer, LOw-Memory Optimization (LOMO), which fuses the gradient computation and the parameter update in one step to reduce memory usage. By integrating LOMO with existing memory saving techniques, we reduce memory usage to 10.8% compared to the standard approach (DeepSpeed solution). Consequently, our approach enables the full parameter fine-tuning of a 65B model on a single machine with 8RTX 3090, each with 24GB memory.111Our code is available at https://github.com/OpenLMLab/LOMO.
@@ 1 Introduction
Recently, parameter-efficient fine-tuning methods (Ding et al., 2022), such as LoRA (Hu et al., 2022) and Prefix-tuning (Li & Liang, 2021), provide solutions for tuning LLMs with limited resources. However, these methods do not offer a practical solution for full parameter fine-tuning, which has been acknowledged as a more powerful approach than parameter-efficient fine-tuning (Ding et al., 2022; Sun et al., 2023). In this work, we aim to explore techniques for accomplishing full parameter fine-tuning in resource-limited scenarios.
@@ 2 Related Work
Mixed-precision training has become a prevalent approach for training large language models due to its ability to accelerate training speed and reduce memory footprint (Narayanan et al., 2021; Rajbhandari et al., 2020). By employing half-precision storage for parameters, activations, and gradients, mixed-precision training enables high-throughput computations during both forward and backward propagation. In order to uphold stability and model accuracy, Micikevicius et al. (2018) proposed three techniques which involve the use of full precision copies of weights, loss scaling, and the execution of specific arithmetic operations in full precision.
@@
Multiple studies (Rhu et al., 2016; Wang et al., 2018; Ren et al., 2021a) have attempted to reduce GPU memory consumption by leveraging heterogeneous memory, such as CPU and NVMe memory. L2L (Pudipeddi et al., 2020) employs a layer-to-layer strategy, where only the tensors necessary for the computation of a particular layer are transferred to the GPU memory, while the remaining tensors are retained in the CPU memory. ZeRO-Offload (Ren et al., 2021b), an extension of ZeRO-2 (Rajbhandari et al., 2020), reserves the gradients and optimizer states in the CPU memory and updates parameters through CPU computation. Tensors and computation operations are allocated to either GPUs or CPUs according to the data-flow graph. ZeRO-Infinity (Rajbhandari et al., 2021), a subsequent advancement of ZeRO-Offload on ZeRO-3 (Rajbhandari et al., 2020), enables further scaling of the model size. Partitioned model states and other tensors can be offloaded not only to CPU memory but also to NVMe to fully leverage heterogeneous architecture.
@@ 3.1 Rethink the Functionality of Optimizer
The optimizer states occupy a large part of the memory used for training LLMs. Modern optimizer like Adam (Kingma & Ba, 2015) stores intermediate states that are twice the size of parameters. As the size of parameters increases, the optimizer states become the dominant term of memory usage.
@@ 3.1.1 Using SGD
Although Adam has achieved great success in training deep models, we ask the question “Can we use a cheaper optimizer for fine-tuning LLMs?” Our answer is SGD, the most basic optimizer. Fortunately, we find that it is an acceptable solution for fine-tuning LLMs when we limit the scope.
Prior works often discuss three challenges of SGD: 1) large curvature loss surface, 2) local optimum, and 3) saddle points (Ruder, 2016; Sun et al., 2020). Modern optimizers have shown effectiveness in dealing with the 1) problem and can mitigate 2) and 3) in some cases. However, when we limit the scope to fine-tuning LLMs, these three challenges could be different.
@@
The goal of fine-tuning is to adapt LLMs to new tasks and domains without significantly changing the model itself. Therefore, a local optimum is often a good enough solution, and the limited training data (compared to pre-training corpus) makes it difficult to push the model to a faraway global optimum.
@@ 3.1.2 Implicit Batch Size
where is a point between and , and we can see that equation 6 minus equation 1 equals the . Suppose the loss surface is smooth enough, this term become negligible. It suggests that utilizing SGD optimizer over a smooth loss surface could imply a larger batch size.
@@ 3.3.1 Alternatives to Gradient Normalization and Clipping
Compute the gradient norm in an additional pass.
@@
Our current training framework computes the gradient norm based on all parameters and requires two backward passes. One solution to save the additional backward pass is to approximate the norm of gradient tensors with a group of parameters, for example, the adjacent layers. However, this method is indeed biased, because it results in different update step sizes for different parameters. When updating, the parameters are multiplied by a scale factor according to the gradient norms. Since the gradient norms differ among parameter groups, such an approximation leads to a difference in scale factors. Despite this limitation, this grouped gradient clipping method can be considered as applying a dynamic learning rate to different groups of parameters based on their gradient norms. Sun et al. (2020) suggests that it is not always appropriate to use the same learning rate for all parameters in SGD, thus we believe our approach also holds the potential to further benefit SGD. We leave the exploration as a compelling future direction.
@@ 4.1 Memory Profile
We first profile the memory usage of model states and activations during the training under different settings. As demonstrated in Table 1, the usage of the LOMO optimizer leads to a substantial reduction in memory footprint from 102.20GB to 14.58GB, when compared to the AdamW optimizer (Loshchilov & Hutter, 2019), and from 51.99GB to 14.58GB, when compared to SGD, in the context of training the LLaMA-7B model. This significant decrease in memory usage can be attributed primarily to the reduced memory requirements of the gradient and optimizer states. As a result, memory is mostly occupied by parameters in the training process, commensurate with memory usage during inference.
@@ 4.3 Downstream Performance
We use the SuperGLUE dataset collection to evaluate model performance, specifically focusing on RTE (Dagan et al., 2005), BoolQ (Clark et al., 2019), WSC (Levesque et al., 2012), WIC (Pilehvar & Camacho-Collados, 2019), MultiRC (Khashabi et al., 2018), and COPA (Roemmele et al., 2011). Given the high computational cost associated with running large language models, we follow MeZO (Malladi et al., 2023) to randomly sample 1000 training data from training set and 1000 test data from validation set, and report the best results obtained using the same random seed. The prompts used in our experiments are the same as MeZO, and the hyperparameters are detailed in Appendix-A.

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) but demand massive GPU resources for training. Lowering the threshold for LLMs training would encourage greater participation from researchers, benefiting both academia and society. While existing approaches have focused on parameter-efficient fine-tuning, which tunes or adds a small number of parameters, few have addressed the challenge of tuning the full parameters of LLMs with limited resources. In this work, we propose a new optimizer, LOw-Memory Optimization (LOMO), which fuses the gradient computation and the parameter update in one step to reduce memory usage. By integrating LOMO with existing memory saving techniques, we reduce memory usage to 10.8% compared to the standard approach (DeepSpeed solution). Consequently, our approach enables the full parameter fine-tuning of a 65B model on a single machine with 8RTX 3090, each with 24GB memory.111Code and data are available at https://github.com/OpenLMLab/LOMO.
Full Parameter Fine-tuning for Large Language Models
with Limited Resources
Kai Lv1,2, Yuqing Yang1, Tengxiao Liu1, Qinghui Gao1, Qipeng Guo2††thanks: Corresponding author., Xipeng Qiu1 1School of Computer Science, Fudan University 2Shanghai AI Laboratory {klv21, yuqingyang21, txliu21}@m.fudan.edu.cn guoqipeng@pjlab.org.cn, xpqiu@fudan.edu.cn
@@ 1 Introduction
Recently, parameter-efficient fine-tuning methods (Ding et al., 2022), such as LoRA (Hu et al., 2022) and Prefix-tuning (Li and Liang, 2021), provide solutions for tuning LLMs with limited resources. However, these methods do not offer a practical solution for full parameter fine-tuning, which has been acknowledged as a more powerful approach than parameter-efficient fine-tuning (Ding et al., 2022; Sun et al., 2023). In this work, we aim to explore techniques for accomplishing full parameter fine-tuning in resource-limited scenarios.
@@ 2 Related Work
Mixed-precision training has become a prevalent approach for training large language models due to its ability to accelerate training speed and reduce memory footprint (Narayanan et al., 2021; Rajbhandari et al., 2020). By employing half-precision storage for parameters, activations, and gradients, mixed-precision training enables high-throughput computations. In order to uphold stability and model accuracy, Micikevicius et al. (2018) proposed three techniques which involve the use of full precision copies of weights, loss scaling, and the execution of specific arithmetic operations in full precision.
@@
Multiple studies (Rhu et al., 2016; Wang et al., 2018; Ren et al., 2021a) have attempted to reduce GPU memory consumption by leveraging heterogeneous memory, such as CPU and NVMe memory. L2L (Pudipeddi et al., 2020) employs a layer-to-layer strategy, where only the tensors necessary for the computation of a particular layer are transferred to the GPU memory, while the remaining tensors are retained in the CPU memory. ZeRO-Offload (Ren et al., 2021b), an extension of ZeRO-2 (Rajbhandari et al., 2020), reserves the gradients and optimizer states in the CPU memory and updates parameters through CPU computation. ZeRO-Infinity (Rajbhandari et al., 2021), a subsequent advancement of ZeRO-Offload on ZeRO-3 (Rajbhandari et al., 2020), enables further scaling of the model size.
In addition to the methods orthogonal to LOMO mentioned above, recent developments have introduced several memory-efficient optimization techniques. MeZO (Malladi et al., 2023) employs a zero-order optimization approach, estimating gradients using two forward passes and updating parameters in place. GaLore (Zhao et al., 2024) performs low-rank decomposition on gradients and uses these approximated gradients for parameter updates. Other approaches reduce memory usage by quantizing optimizer states (Dettmers et al., 2022; Sun et al., 2020b). Compared to these methods, LOMO neither approximates gradients nor requires low-bit quantization.
@@ 3.1 Rethink the Functionality of Optimizer
The optimizer states occupy a large part of the memory used for training LLMs. Modern optimizer like Adam (Kingma and Ba, 2015) stores intermediate states that are twice the size of parameters. As the size of parameters increases, the optimizer states become the dominant term of memory usage.
@@ 3.1.1 Using SGD
Although Adam has achieved great success in training deep models, we ask the question “Can we use a cheaper optimizer for fine-tuning LLMs?" Our answer is SGD, the most basic optimizer. Fortunately, we find that it is an acceptable solution for fine-tuning LLMs when we limit the scope.
Prior works often discuss three challenges of SGD: 1) large curvature loss surface, 2) local optimum, and 3) saddle points (Ruder, 2016; Sun et al., 2020a). Modern optimizers have shown effectiveness in dealing with the 1) problem and can mitigate 2) and 3) in some cases. However, when we limit the scope to fine-tuning LLMs, these three challenges could be different.
@@
The goal of fine-tuning is to adapt LLMs to new tasks and domains without significantly changing the model itself. Therefore, a local optimum is often a good enough solution Kawaguchi et al. (2019), and the limited training data (compared to pre-training corpus) makes it difficult to push the model to a faraway global optimum.
@@ 3.1.2 Implicit Batch Size
where is a point between and , and we can see that (6) minus (1) equals the . Suppose the loss surface is smooth enough, this term become negligible. It suggests that utilizing SGD optimizer over a smooth loss surface could imply a larger batch size.
@@ 3.3.1 Alternatives to Gradient Normalization and Clipping
Compute the gradient norm in an additional backward pass.
@@
Our current training framework computes the gradient norm based on all parameters and requires two backward passes. One solution to save the additional backward pass is to approximate the norm of gradient tensors with a group of parameters, for example, the adjacent layers. This method is indeed biased, because it results in different update step sizes for different parameters. When updating, the parameters are multiplied by a scale factor according to the gradient norms. Since the gradient norms differ among parameter groups, such an approximation leads to a difference in scale factors. This grouped gradient clipping method can be considered as applying a dynamic learning rate to different groups of parameters based on their gradient norms. Sun et al. (2020a) suggests that it is not always appropriate to use the same learning rate for all parameters in SGD, thus we believe our approach also holds the potential to further benefit SGD. We leave the exploration as a compelling future direction.
@@ 3.3.2 Mitigating Precision Degradation
4 Experiment
4.1 Memory Profile
We first profile the memory usage of model states and activations during the training under different settings. As demonstrated in Table 1, the usage of the LOMO optimizer leads to a substantial reduction in memory footprint from 102.20GB to 14.58GB, when compared to the AdamW optimizer (Loshchilov and Hutter, 2019), and from 51.99GB to 14.58GB, when compared to SGD, in the context of training the LLaMA-7B model. This significant decrease in memory usage can be attributed primarily to the reduced memory requirements of the gradient and optimizer states. As a result, memory is mostly occupied by parameters in the training process, commensurate with memory usage during inference.
@@ 4.3 Downstream Performance
We use the SuperGLUE dataset collection to evaluate model performance, specifically focusing on RTE (Dagan et al., 2005), BoolQ (Clark et al., 2019), WSC (Levesque et al., 2012), WIC (Pilehvar and Camacho-Collados, 2019), MultiRC (Khashabi et al., 2018), and COPA (Roemmele et al., 2011). Given the high computational cost associated with running large language models, we follow MeZO (Malladi et al., 2023) to randomly sample 1000 training data from training set and 1000 test data from validation set, and report the best results obtained using the same random seed. The prompts used in our experiments are the same as MeZO, and the hyperparameters are detailed in Appendix-A.
@@ 5 Conclusion
Limitations
In response to the challenges associated with gradient normalization and clipping, we have developed alternative optimization methods. Although gradient normalization for LOMO does not increase memory usage, our current implementation necessitates an additional backward pass, which can slow down the training speed in scenarios where gradient normalization is essential.
Due to time and resource constraints, our experiments were limited to a subset of the SuperGLUE benchmark, and we did not evaluate LOMO’s throughput on advanced GPUs such as A100.
Ethics statement
This paper employs open-source models LLaMA, in compliance with their respective licenses. The datasets utilized, including RTE, BoolQ, WSC, WIC, MultiRC and COPA, permit public and freeusage.
Acknowledgments
This work was supported by the National Key Research and Development Program of China (No.2022ZD0160102). The computations in this research were performed using the CFFF platform of Fudan University.
@@ Appendix A Hyperparameters
Appendix B Training Dynamics
To analyze the training dynamics of LOMO, we present the training loss curve and validation accuracy for LLaMA-7B trained on BoolQ (Clark et al., 2019) using LOMO and LoRA in Figure 4 and Figure 5, respectively. During training process with LOMO, the loss converges rapidly in the initial phase and then tends to stabilize and gradually decline. The accuracy on the development set generally shows an upward trend as the number of training steps increases.